OpenStudy (anonymous):

At a local bank, a teller noticed that for every 17 checks that customers deposit with the bank, 5 take fewer than three days to process completely. Last week, 150 checks took fewer than three days to process completely. What is a reasonable estimate of the total number of checks that customers deposited with the bank last week?

OpenStudy (anonymous):

5/17=x/150 x=(150x5)/17 x=about 44

ganeshie8 (ganeshie8):

\(\large \frac{17}{5} = \frac{x}{150}\)

OpenStudy (anonymous):

470 510 450 550

OpenStudy (anonymous):

oh i set that up wrong hold on pls

OpenStudy (anonymous):

using @ganeshie8 's equation: x=(150x17)/5 x=510